e.com/platform/mod/quiz/attempt.php?attempt-23468068cid 1691668page 58scrollpos-440 nu Equivalent Units Calculations-FIFO Method Ferris Corporation makes a powdered rug shampoo in two sequential departments, Compounding and Drying. Materials are added at the beginning of the process in the Compounding Department. Conversion costs are added evenly throughout each process. Ferris uses the FIFO method of process costing. In the Compounding Department, beginning work in process was 14,000 pounds (65% processed), 129,500 pounds were started in process, 126,000 pounds transferred out, and ending work in process was 65% processed. Calculate equivalent units for the Compounding Department for March 2016. Ferris Corporation Flow of Units and Equivalent Units Calculation, March 2016 Equivalent Units Direct MarchMaterialsMarch %in Conversion Costs Complete/Transferred Beginning inventory 14,000 35 4,900 Started & 8,400 100 100 completec 9,600 100 65 Ending inventory Total
Answer #1

Under FIFO ,units started first are transferred first so out of the units transferred 126000 ,14000 is from beginning WIP and remaining is transferred from units started and completed during the year [126000-14000= 112000]

Ending WIP = 14000+129500-126000=17500

Ferris corporation

Flow of units and equivalent unit calculation ,march 2016

Equivalent unit
% in march Direct material % of March conversion cost
completed and transferred
Beginning Inventory 14000 0% 0 35% 4900
started and completed 112000 100% 112000 100% 112000
Ending inventory 17500 100% 17500 65% 11375
Total 143500 129500 128275